1 0
Add a comment Improve this question Transcribed image text
Answer #1
Concepts and reason

This problem is based on the concept of conversion of potential energy to force on atom and equilibrium distance between two atoms.

Initially, differentiate the potential energy expression to calculate the expression for the force with a negative sign, calculate the equilibrium distance between atoms.

Later, if potential energy is written in an expression, the differential of the potential energy with negative sign gives the force on the atom.

Finally, equilibrium distance is measured by using the expression of the force on the atom, this expression is equaled to zero to find the equilibrium distance.

Fundamentals

The force F(r)F(r) is written as follows:

F(r)=ddr[U(r)]F(r) = - \frac{d}{{dr}}\left[ {U(r)} \right]

Here, U(r)U(r) is the function of potential energy.

The force F(r)F(r) should be zero, therefore it is written as follows:

F(r)=[12ar136br7]=0\begin{array}{c}\\F(r) = \left[ {\frac{{12a}}{{{r^{13}}}} - \frac{{6b}}{{{r^7}}}} \right]\\\\ = 0\\\end{array}

Here, rr is the spacing between atoms, aa and bb are positive constants.

The case for stability is given as,

dF(r)dr<0\frac{{dF(r)}}{{dr}} < 0

Here, F(r)F(r) is the function of force.

(a)

The expression for the potential energy of two atoms in a diatomic molecule is approximated as,

U(r)=ar12br6U(r) = \frac{a}{{{r^{12}}}} - \frac{b}{{{r^6}}} …… (1)

The force F(r)F(r) is written as,

F(r)=ddr[U(r)]F(r) = - \frac{d}{{dr}}\left[ {U(r)} \right] …… (2)

From expression (1) and (2),

F(r)=ddr[ar12br6]=[12ar136br7]=[12ar136br7]\begin{array}{c}\\F(r) = - \frac{d}{{dr}}\left[ {\frac{a}{{{r^{12}}}} - \frac{b}{{{r^6}}}} \right]\\\\ = - \left[ {\frac{{ - 12a}}{{{r^{13}}}} - \frac{{ - 6b}}{{{r^7}}}} \right]\\\\ = \left[ {\frac{{12a}}{{{r^{13}}}} - \frac{{6b}}{{{r^7}}}} \right]\\\end{array}

(b)

The equilibrium distance between two atoms is calculated as follows:

F(r)=[12ar136br7]=0\begin{array}{c}\\F(r) = \left[ {\frac{{12a}}{{{r^{13}}}} - \frac{{6b}}{{{r^7}}}} \right]\\\\ = 0\\\end{array}

Further, it is written as follows:

[12ar136br7]=012ar13=6br7r6=2ab\begin{array}{l}\\\left[ {\frac{{12a}}{{{r^{13}}}} - \frac{{6b}}{{{r^7}}}} \right] = 0\\\\\frac{{12a}}{{{r^{13}}}} = \frac{{6b}}{{{r^7}}}\\\\{r^6} = \frac{{2a}}{b}\\\end{array}

Solve for r.

r=(2ab)16r = {\left( {\frac{{2a}}{b}} \right)^{\frac{1}{6}}}

(c)

The case for stability is given as,

dF(r)dr<0\frac{{dF(r)}}{{dr}} < 0

The differentiation of F(r)F(r) with respect to rr is written as follows:

dF(r)dr=ddr[12ar136br7]=[(12)(13)ar14+(6)(7)br8]=[156ar14+42br8]\begin{array}{c}\\\frac{{dF(r)}}{{dr}} = \frac{d}{{dr}}\left[ {\frac{{12a}}{{{r^{13}}}} - \frac{{6b}}{{{r^7}}}} \right]\\\\ = \left[ {\frac{{ - \left( {12} \right)\left( {13} \right)a}}{{{r^{14}}}} + \frac{{\left( 6 \right)\left( 7 \right)b}}{{{r^8}}}} \right]\\\\ = \left[ {\frac{{ - 156a}}{{{r^{14}}}} + \frac{{42b}}{{{r^8}}}} \right]\\\end{array}

Substitute the value (2ab)16{\left( {\frac{{2a}}{b}} \right)^{\frac{1}{6}}} for rr in above expression.

dF(r)dr=[156a((2ab)16)14+42b((2ab)16)8]=[156a(2ab)146+42b(2ab)86]<0\begin{array}{c}\\\frac{{dF(r)}}{{dr}} = \left[ {\frac{{ - 156a}}{{{{\left( {{{\left( {\frac{{2a}}{b}} \right)}^{\frac{1}{6}}}} \right)}^{14}}}} + \frac{{42b}}{{{{\left( {{{\left( {\frac{{2a}}{b}} \right)}^{\frac{1}{6}}}} \right)}^8}}}} \right]\\\\ = \left[ {\frac{{ - 156a}}{{{{\left( {\frac{{2a}}{b}} \right)}^{\frac{{14}}{6}}}}} + \frac{{42b}}{{{{\left( {\frac{{2a}}{b}} \right)}^{\frac{8}{6}}}}}} \right]\\\\ < 0\\\end{array}

(d)

Substitute a(2ab)2- a{\left( {\frac{{2a}}{b}} \right)^{ - 2}} for Ufinal{U_{{\rm{final}}}} and b(2ab)1b{\left( {\frac{{2a}}{b}} \right)^{ - 1}} for Uinitial{U_{{\rm{initial}}}} in the equation Uadded=UfinalUinitial{U_{{\rm{added}}}} = {U_{{\rm{final}}}} - {U_{{\rm{initial}}}} .

Uadded=a(2ab)2b(2ab)1=b24a+b22a=b24a\begin{array}{c}\\{U_{{\rm{added}}}} = - a{\left( {\frac{{2a}}{b}} \right)^{ - 2}} - b{\left( {\frac{{2a}}{b}} \right)^{ - 1}}\\\\ = - \frac{{{b^2}}}{{4a}} + \frac{{{b^2}}}{{2a}}\\\\ = \frac{{{b^2}}}{{4a}}\\\end{array}

(e)

The value of the aquarium distance for the Co molecule is equal to,

r0=(2ab)16=1.13×1010m\begin{array}{c}\\{r_0} = {\left( {\frac{{2a}}{b}} \right)^{\frac{1}{6}}}\\\\ = 1.13 \times {10^{ - 10}}{\rm{ m}}\\\end{array}

And the dissociation energy is as follows:

Uadded=b24a=1.54×1018\begin{array}{c}\\{U_{{\rm{added}}}} = \frac{{{b^2}}}{{4a}}\\\\ = 1.54 \times {10^{ - 18}}{\rm{ }}\\\end{array}

Therefore, from the equation r0=(2ab)16{r_0} = {\left( {\frac{{2a}}{b}} \right)^{\frac{1}{6}}} and b24a\frac{{{b^2}}}{{4a}} ,

2ab=2.08×1060\frac{{2a}}{b} = 2.08 \times {10^{ - 60}}

Using the equation b24a=1.54×1018m\frac{{{b^2}}}{{4a}} = 1.54 \times {10^{ - 18}}{\rm{ m}} .

b2=4a(1.54×1018m){b^2} = 4a\left( {1.54 \times {{10}^{ - 18}}{\rm{ m}}} \right)

Solve for a.

a=6.656×10138a = 6.656 \times {10^{ - 138}}

(f)

Using the equation b24a=1.54×1018m\frac{{{b^2}}}{{4a}} = 1.54 \times {10^{ - 18}}{\rm{ m}} .

b2=4a(1.54×1018m){b^2} = 4a\left( {1.54 \times {{10}^{ - 18}}{\rm{ m}}} \right)

Substitute 6.656×101386.656 \times {10^{ - 138}} for a.

b2=4(6.656×10138)(1.54×1018m)b=4(6.656×10138)(1.54×1018m)=6.40×1078\begin{array}{c}\\{b^2} = 4\left( {6.656 \times {{10}^{ - 138}}} \right)\left( {1.54 \times {{10}^{ - 18}}{\rm{ m}}} \right)\\\\b = \sqrt {4\left( {6.656 \times {{10}^{ - 138}}} \right)\left( {1.54 \times {{10}^{ - 18}}{\rm{ m}}} \right)} \\\\ = 6.40 \times {10^{ - 78}}\\\end{array}

Ans: Part a

The force F(r)F(r) on one atom as a function of rr is [12ar136br7]\left[ {\frac{{12a}}{{{r^{13}}}} - \frac{{6b}}{{{r^7}}}} \right] .

Part b

The equilibrium distance between the two atoms is (2ab)16{\left( {\frac{{2a}}{b}} \right)^{\frac{1}{6}}} .

Part c

This Equilibrium is stable.

Part d

The dissociation energy is equal to b24a\frac{{{b^2}}}{{4a}} .

Part e

The value of the constant a is equal to 6.656×101386.656 \times {10^{ - 138}} .

Part f

The value of the constant b is equal to 6.40×10786.40 \times {10^{ - 78}} .